# Loss Functions for Measuring the Accuracy of Nonnegative Cross-Sectional Predictions

### v1

@@ 1 Introduction
Section 2 axiomatically constructs the simple loss function that we will use. It further specifies assumptions and develops the simplest loss function that satisfies those assumptions also has the property that the loss associated with a given absolute relative error [^7] Equivalently, one can also refer to the absolute percentage error (APE), but this will complicate the arguments made later on. APE will appear, however, in the example in Section 6. increases in the actual value. Section 3 discusses estimation of this loss function, along with specification tests. Section 4 argues for the use of the Webster-Saint Lagüe Rule when predictions can be viewed as apportionments. Given a set of predicted and actual values, the total loss function is the sum of the loss functions for the individual observations. Different sets of predictions may have systematic, offsetting biases. The decision-maker may wish to form weighted averages of these predictions to produce a set of new predictions. Subection 3.4 provides a method to search for the optimal weightings which minimize the total loss function. Section 5 provides an example of the use of loss functions, comparing the Webster-Saint Lagüe Rule to the commonly used mean absolute percentage error (MAPE).
@@ 3.2.1 Multiple Decision Makers
The fundamental problem is that there is no way to specify a unique utility function with which to evaluate prediction errors. Arrow’s (1950) famous theorem disproving the existence of a function that aggregates individuals’ preferences and obeys some weak conditions underlies the last claim.
@@ 4 Predictions as Apportionments: The Webster-Saint Lagüe Rule
Often predictions represent apportionments, that is, resource or legislative seat allocations. The U.S. decennial census is used to apportion the House of Representatives. Population estimates are used by some states to apportion tax revenues to localities. One would like to have the rule used for evaluating the predictions satisfy various fairness criteria. Balinski and Young (1982) have shown that the Webster-Saint Lagüe Rule satisfies a large number of fairness criteria. [^29] However, no method can simultaneously satisfy all fairness criteria (Balinski and Young, 1979, th. 6). For example, Ernst (1992) proves that Hill’s method satisfies other fairness criteria. This result is similar to Arrow’s (1950) Theorem, mentioned above in Subsection 3.2.1. Still, the Webster-Saint Lagüe Rule is the only apportionment rule considered by Balinski and Young (1979, 1982) and Ernst (1992) that can be represented by minimizing the sum of loss functions satisifying Assumptions 1-3 and Property 1 (Spencer, 1985). I would like to thank Gregg Diffendal for informing me of Ernst (1992). Most of these criteria result from the Webster-Saint Lagüe Rule being a divisor method. the Webster-Saint Lagüe Rule method is the only divisor method which has an admissible (in the present paper) loss function representation: $p=2$ and $q=-1$, (Balinski and Young, 1985) See Subsection 6.5 below for other apportionment-based loss functions. An additional normative criterion of particular interest is unbiasedness: for any two disjoint groups of areas, with all members in one group larger than those in the other, the probability that the method favors large areas equals the probability that it favors the small areas. (Balinksi and Young, 2001, Theorem 5.3)
@@ 4.2 Fellegi’s Variant
Fellegi (1980, eq. 8, 194) proposes an inequality measure based on differences between optimal and predicted shares. He uses conventions to assign exponents. Hogan and Mulry (2014, eq. 94, 124) show that Fellegi’s measure is equivalent to the loss function [^30] This is after removing constants and the summation.
@@ 5 Example of Evaluating Predictions Using Loss Functions
Loss functions can produce entirely different and more meaningful results than common error measures such as the mean absolute percentage error (MAPE). [^31] That is the loss function is computed with $p=1$ and $q=-1$, absolute values taken, then multiplied by 100 for reexpression in terms of percentage points. Table 1 shows the true values of six areas, $A_{i},i=1,\dots,6$, and three sets (Scenarios) of absolute errors $(\epsilon_{i})$, along with the corresponding absolute percentage errors ($\mathrm{APE}_{i})$ and Webster-Lagüe Rule loss function values $L_{i}$. [^32] This example is from Coleman (2000). The bottom row shows the means of the last two variables. These are simply MAPE and $\mathcal{L}_{W}/n$, respectively.
@@ 6.2 Median Absolute Percentage Error
Smith and Sincich (1992) use the 90th Percentile Absolute Percentage Error (90PE) as one of their metrics for comparing cross-sectional forecasts. Like MedAPE, this can only be defined in our framework by the total loss function, $100*\mathrm{median}_{i}\lvert\frac{P_{90}-A_{90}}{A_{90}}\rvert$, where the subscript 90 refers to the observation whose absolute percentage error is greater than that of 90% of all the observations. 90PE shares MedAPE’s weaknesses, while being sensitive to fewer extreme errors.
@@ 6.4 Root Mean Squared Percentage Error
The Root Mean Squared Percentage Error (RMSPE) is defined as $\sqrt{\frac{[\sum_{i=1}^{n}(P_{i}-A_{i})/{\sum_{i=1}^{n}A_{i}}]^{2}}{n}}$. This is equivalent to $p=2$ and $q=-2$. Like MAPE, this violates Property 1. It is useful only when the a priori relative mean squared errors are constant. That is, $\mathbf{E}(\frac{P_{i}-A_{i}}{A_{i}})=c$, for some constant $c>0$.
@@ 6.5 Apportionment-Based Loss Functions
Spencer (1985) develops five apportionment-based loss functions. In addition to the Webster-Saint Lagüe Rule, these are (1) $p=1$, $q=0$, or more generally, $p>0,q=0$(Hamilton’s method), (2) $p=2,q=-1$ with $P_{i}$ raised to the $q$ instead of $A_{i}$. (Hill’s method/Method of Equal Propertions), (3) $\max_{i}P_{i}/A_{i}$ (Jefferson/d’Hondt method) and (4) $\max_{i}A_{i}/P_{i}$ (Adam’s Method). Hamilton’s and Hill’s methods both violate Assumption 3. The Jefferson/d’Hondt method violates both Assumption 3 and the von Neumann-Morgenstern axioms. Adam’s Method additionally violates Assumptions 1 and 2. Spencer (1986, p. 398) advocates incorporating expectations into these methods to measure accuracy. However, since we have assumed the distributions generating the errors to be unknown, this is impossible.
@@ 7 Loss Function-Based Bias Measures
Any of our loss functions can be converted to a bias measure simply by multiplying loss by the sign of $\epsilon$. From page 14 we know that $\epsilon=|P-A|$. Therefore, equation (1b) becomes

### v2

@@ 1 Introduction
Section 2 axiomatically constructs the simple loss function that we will use. It further specifies assumptions and develops the simplest loss function that satisfies those assumptions also has the property that the loss associated with a given absolute relative error [^7] Equivalently, one can also refer to the absolute percentage error (APE), but this will complicate the arguments made later on. APE will appear, however, in the example in Section 6. increases in the actual value. Section 3 discusses estimation of this loss function, along with specification tests. Different sets of predictions may have systematic, offsetting biases. The decision-maker may wish to form weighted averages of these predictions to produce a set of new predictions. Subection 3.4 provides a method to search for the optimal weightings which minimize the total loss function. Section 4 argues for the use of the Webster-Saint Lagüe Rule when predictions can be viewed as apportionments. Given a set of predicted and actual values, the total loss function is the sum of the loss functions for the individual observations. Section 5 provides an example of the use of loss functions, comparing the Webster-Saint Lagüe Rule to the commonly used mean absolute percentage error (MAPE).
@@ 3.2.1 Multiple Decision Makers
The fundamental problem is that there is no way to specify a unique utility function with which to evaluate prediction errors. Arrow’s (1950) famous Impossibility Theorem disproving the existence of a function that aggregates individuals’ preferences and obeys some weak conditions underlies the last claim.
@@ 4 Predictions as Apportionments: The Webster-Saint Lagüe Rule
Often predictions represent apportionments, that is, resource or legislative seat allocations. The U.S. decennial census is used to apportion the House of Representatives. Population estimates are used by some states to apportion tax revenues to localities. One would like to have the rule used for evaluating the predictions satisfy various fairness criteria. Balinski and Young (2001) show that the Webster-Saint Lagüe Rule satisfies a large number of fairness criteria. [^29] However, while Arrow’s( 1950) Impossibility Theorem implies that no voting system can simultaneously satisfy all fairness criteria, the Webster-Saint Lagüe Rule is the only apportionment rule considered by Balinski and Young (1979, 2001) and Ernst (1992) that can be represented by minimizing the sum of loss functions satisifying Assumptions 1-3 and Property 1 (Spencer, 1985). I would like to thank Gregg Diffendal for informing me of Ernst (1992). Most of these criteria result from the Webster-Saint Lagüe Rule being a divisor method. the Webster-Saint Lagüe Rule method is the only divisor method which has an admissible (in the present paper) loss function representation: $p=2$ and $q=-1$, (Balinski and Young, 1985) See Subsection 6.5 below for other apportionment-based loss functions. An additional normative criterion of particular interest is unbiasedness: for any two disjoint groups of areas, with all members in one group larger than those in the other, the probability that the method favors large areas equals the probability that it favors the small areas. (Balinksi and Young, 2001, Theorem 5.3)
@@ 4.2 Fellegi’s Variant
Fellegi (1980, eq. 8, 194) proposes an inequality measure based on differences between optimal and predicted shares. He uses conventions to assign exponents. Hogan and Mulry (2014, eq. 94, p. 124) show that Fellegi’s measure is equivalent to the loss function [^30] This is after removing constants and the summation.
@@ 5 Example of Evaluating Predictions Using Loss Functions
Loss functions can produce entirely different and more meaningful results than common error measures such as the mean absolute percentage error (MAPE). [^31] That is the loss function is computed with $p=1$ and $q=-1$, absolute values taken, then multiplied by 100 for reexpression in terms of percentage points. Table 1 shows the true values of six areas, $A_{i},i=1,\dots,6$, and three sets (Scenarios) of absolute errors $(\epsilon_{i})$, along with the corresponding absolute percentage errors ($\mathrm{APE}_{i})$ and Webster-Saint Lagüe Rule loss function values $L_{i}$. [^32] This example is from Coleman (2000). The bottom row shows the means of the last two variables. These are simply MAPE and $\mathcal{L}_{W}/n$, respectively.
@@ 6.2 Median Absolute Percentage Error
Smith and Sincich (1992) use the 90th Percentile Absolute Percentage Error (90PE) as one of their metrics for comparing cross-sectional forecasts. Like MedAPE, this can only be defined in our framework by the total loss function, $100*\lvert\frac{P_{90}-A_{90}}{A_{90}}\rvert$, where the subscript 90 refers to the observation whose absolute percentage error is greater than that of 90% of all the observations. 90PE shares MedAPE’s weaknesses, while being sensitive to fewer extreme errors.
@@ 6.4 Root Mean Squared Percentage Error
The Root Mean Squared Percentage Error (RMSPE) is defined as $\sqrt{\frac{[\sum_{i=1}^{n}(P_{i}-A_{i})/{\sum_{i=1}^{n}A_{i}}]^{2}}{n}}$. This is equivalent to $p=2$ and $q=-2$. Like MAPE, this violates Property 1. It is useful only when the a priori relative mean squared errors are constant. That is, $\mathbf{E}(\frac{P_{i}-A_{i}}{A_{i}})^{2}=c$, for some constant $c>0$.
@@ 6.5 Apportionment-Based Loss Functions
Spencer (1985) develops five apportionment-based loss functions. In addition to the Webster-Saint Lagüe Rule, these are (1) $p=1$, $q=0$, or more generally, $p>0,q=0$. (Hamilton’s method), (2) $p=2,q=-1$ with $P_{i}$ raised to the $q$ instead of $A_{i}$ (Hill’s method/Method of Equal Propertions), (3) $\max_{i}P_{i}/A_{i}$ (Jefferson/d’Hondt method) and (4) $\max_{i}A_{i}/P_{i}$ (Adam’s Method). Hamilton’s and Hill’s methods both violate Assumption 3. The Jefferson/d’Hondt method violates both Assumption 3 and the von Neumann-Morgenstern axioms. Adam’s Method additionally violates Assumptions 1 and 2. Spencer (1986, p. 398) advocates incorporating expectations into these methods to measure accuracy. However, since we have assumed the distributions generating the errors to be unknown, this is impossible.
@@ 7 Loss Function-Based Bias Measures
Any of our loss functions can be converted to a bias measure simply by multiplying loss by the sign of $\epsilon$, similar to the construction of the Mean Algebraic Percentage Error (MALPE) from MAPE. From page 14 we know that $\epsilon=|P-A|$. Therefore, equation (1b) becomes